# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

### v2

@@ 6.2 Verification against radiosondes
Comparison of forecasts at full resolution against the global radiosonde observations clearly indicates that GDPS-SN forecasts are closer to the observations than GDPS-CTL forecasts for basically all variables, lead times and levels, and over both seasons (Fig. 9). The improvements gradually grow in the first 72 hr and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GC in Section 4. This demonstrates that large-scale spectral nudging, despite being only applied in the mid-level layer for winds and temperature, is sufficient to impart similar improvements at both upper- and lower-level layers that are not directly constrained by spectral nudging. Compared to boreal winter (Fig. 9a), the improvements from GDPS-SN are reduced by a factor of about 2 during boreal summer (Fig. 9b), which is not surprising given that the large-scale improvements with GC over GDPS-CTL are smaller by about the same factor during boreal summer (not shown).
@@
In terms of region-wise performances, the GDPS-SN has smaller RMSE than GDPS-CTL in every subdomain except over Asia in the mid- and upper-levels during boreal summer (Figs. 10c and d), which is consistent with the degradations from GC over the GDPS-CTL at the large scales in this region (Fig. 5d). Unlike GC, the improvements from GDPS-SN over GDPS-CTL are virtually identical when only the large-scale components of the forecast are compared to the radiosonde observations (compare the left and right columns in Fig. 10 and, again, in Fig. 5). This does not necessarily imply that the improvements only originate from the large scales. It rather implies that the relative RMSE reductions with GDPS-SN are of similar level for both the large and small scales. In other words, as the small scales are generally conditioned by the large scales, these results suggest that the fine scales predicted by GDPS-SN benefit from fractional improvements that are on par with its large-scale skill enhancement.
@@
Figure 9: Same as in Fig. 4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite. Only verification using forecasts at full resolution is shown.
@@
Figure 10: Same as in Fig. 5, except that red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
@@
Figure 11: Same as in Fig. 4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to GC, whereas downward-pointing blue triangles indicate the opposite. Only verification using forecasts at low resolution is shown but for both seasons.
@@
It is also interesting to compare GDPS-SN to GC. For the sake of fair comparison, this is only done at low resolution. Predictions from GDPS-SN are closer to the radiosonde observations than GC in the high- and mid-level layers, except for dewpoint depression. Conversely, for the low-level layer, GC is closer to the observations, except for geopotential height (Fig. 11). The degradations with respect to dewpoint depression, also observed in the mid-level layer where spectral nudging is active, suggest that further gains could likely be achieved by nudging the humidity field in GEM towards GC. Nudging of specific humidity, however, would only be viable once a fine-tuned version of GC that emulates GDPS analyses instead of ERA5 (and ECMWF-HRES) is available. This is expected to bring the mean moisture state of GC inferences closer to GDPS analyses and thereby reduce any error associated with potential adverse response from GEM physics. The results also suggest that extending spectral nudging to the boundary layer could be beneficial. This, however, would necessitate an AI model with more vertical levels in the boundary layer (as mentioned earlier) and fine-tuned to emulate the GDPS analyses rather than ECMWF-ERA5, for improved consistency with the GEM model’s lower-boundary forcings.
@@
The small but statistically significant GDPS-SN improvements over GC in the mid-level layer for temperature and zonal winds, two variables that are nudged toward GC, that grow with the integration time, are somewhat surprising (Fig. 11). The use of a nudging relaxation time $\tau=12$ hr implies that the large-scales in GDPS-SN are not too tightly constrained, which is visible in Fig. 8. This likely allows for some large-scale improvements through upscale propagation of information from an improved representation of the fine scales. However, this choice of $\tau$ also leads to a small averaging effect caused by the blending of two forecasts, as previously discussed in Section 5. To a lesser extent, this may also have contributed to these improvements.
@@
Finally, the overall region-wise comparison of GDPS-SN and GC forecasts at low resolution (Fig. 12) shows that the skill of GDPS-SN is either slightly better than or comparable to GC, with the tropical boundary layer being an exception. Fig. 5 reveals that, compared to GC, GDPS-CTL large scales have substantially reduced skill in the tropics, which is slightly improved with GDPS-SN (see Fig. 10). To leverage the full potential of GC in the tropics it appears to be necessary to extend nudging in the boundary layer, which is not viable at this stage for reasons explained earlier in this section.
@@
Figure 12: Same as in Fig. 5, except that red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GC, whereas blue shadings indicate the opposite. Only verification using forecasts at low resolution is shown.
@@ 6.3 Verification against ECMWF analyses
The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF-HRES analyses leads to very similar results as the comparison against radiosonde observations, with general improvements for all variables, lead times and pressure levels (not shown). Figure 13 provides an example of improvements in anomaly correlation coefficient (ACC). For 500-hPa geopotential heights, the ACC improvements over the Northern Hemisphere region at forecast day 7 are equivalent to an 13-hr increase in predictability during boreal winter and to an 8-hr increase in summer. The corresponding ACC improvements over the Southern Hemisphere region are about 3 hours during austral winter and 22 hours during austral summer. There are also notable improvements in zonal wind ACC in the tropics that are equivalent to 30 hours or more (e.g., see Fig. 13b,e). The strength of GC in improving the ACC around day 5 (and beyond) is thus well leveraged by the hybrid system. Note that GC has not been included in this figure as its ACC is strongly influenced by the excessive fine-scale smoothing. To fairly compare GC against the two versions of GDPS, it is necessary to do that at low resolution (as in section 4b). However, based on spectral coherence presented in Fig. 7c–d, low-resolution ACC of GC is expected to be comparable to GDPS-SN.
@@
Figure 13: Anomaly correlation coefficient against the ECMWF-HRES analyses with GDPS-CTL (blue) and GDPS-SN (red) for (a,d) northern hemisphere 500-hPa geopotential height, (b,e) tropical 850-hPa zonal wind, and (c,f) southern hemisphere 500-hPa geopotential height. Results are presented for 60 cases of boreal winter 2022 (left column: a–c) and 62 cases of boreal summer 2022 (right column: d–f). Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 95th percentile. Grey arrows with printed numbers depict gains in predictability in terms of forecast hours.
6.4 Verification against surface observations
@@
For verification at the surface level, combined SYNOP, METAR and Surface Weather and marine OBservations (SWOB; available only over Canada and distributed by ECCC) data were used for surface pressure, screen-level temperature, dewpoint depression and anemometer-level wind speed. For precipitation, ground observations of 24-hr accumulation, subject to collection and quality control by the CAnadian Precipitation Analysis system (CaPA; Lespinas et al. 2015), are used. Any data pertaining to stations with an altitude difference larger than 100 m with respect to GDPS orography, along with any observed wind speed below 1.5 $\text{m s}^{-1}$, are excluded.
@@
Figure 14: Heatmap of the changes in the forecast quality index (Eq. 3) against combined SYNOP, SWOB and METAR observations for surface pressure ($P_{s}$), screen-level temperature ($T$), dewpoint depression ($T-T_{d}$) and anemometer-level wind speed ($W$) in various geographical domains for (a) 60 cases of boreal winter 2022 and (b) 62 cases of boreal summer 2022. Red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
@@
Figure 15: Q-Q plots of day-5 GC (black), GDPS-CTL (blue) and GDPS-SN (red) forecasts against combined SYNOP, SWOB and METAR for: (a, b) anemometer-level wind speeds and (c, d) screen-level temperatures over North America for (a, c) 30 cases of boreal winter valid at (a) 00 and (c) 12 UTC, and (b, d) 31 cases of boreal summer valid 00 UTC. Local standard time over North America is mostly between UTC-5 and UTC-8 h. Diamond symbols denote selected percentiles with their values printed. Light shadings indicate the 5%–95% confidence interval for the inverse cumulative distribution functions, based on the Kolmogorov-Smirnov statistic.
@@
Comparison of GDPS-SN and GDPS-CTL using the forecast quality index (Eq. 3) is summarized in Fig. 14. Guidance by GDPS-SN is generally improved for all variables and regions. The improvements are, however, notably smaller than in the upper air, which is particularly the case for the tropics, where achieving considerable improvements appears to be difficult without introducing spectral nudging in the boundary layer. The same could be concluded for the screen-level thermodynamic variables and wind. On the other hand, the spectral nudging in the mid-level layer has the largest positive impact on the surface pressure via redistribution of upper-air mass. The improvements in boreal summer are notably smaller than in boreal winter, consistent with the upper-air verification.
@@
The ability of the GPDS-SN to represent high-impact weather, such as high wind speeds and extreme temperatures, is not significantly different from the GDPS-CTL but, at the same time, it is substantially better than GC, which provides a strong example of the value added by the hybrid system over GC (Fig. 15). For example, the distribution of late-afternoon wind speeds over North America provided by GC shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -3 $\text{m s}^{-1}$ bias around ${}^{\text{th}}$ percentile (Figs. 15 a, b). GDPS-CTL also has a negative bias but affecting only high percentiles and, more importantly, GDPS-SN closely follows GDPS-CTL, with little or no sign of deceleration at extreme wind speeds for both seasons. GraphCast also suffers a 2-4 K cold bias at mid-to-high temperature percentiles over North America in summer, which has no impact on the GDPS-SN; its distribution aligns very well with the observations. GraphCast’s poor performance with respect to the extremes can be partly explained by its lack of well-resolved fine scales that are generally critical for representing the extremes. Furthermore, inconsistencies in the surface forcing between ERA5 and GDPS analyses may also be a contributing factor for the observed underestimation of temperature extremes by GC. As the spectral nudging configuration presented in this study only targets synoptic scales and the boundary layer is allowed to evolve freely, GDPS-SN is not adversely affected by GC’s limitations regarding the extremes.
@@
Spectral nudging also significantly improves boreal winter precipitation forecast guidance over North America for a large range of thresholds from 1 to 50 mm day-1, mostly by reducing the false alarm ratio, with no reduction of probability of detection, which results in a reduced moist frequency bias and an improved equitable threat score (Fig. 16). These improvements seem to be mostly generated via an improved prediction of synoptic-scale flow. There is, however, not much of an impact of spectral nudging on summer precipitation over North America (not shown). During summer, when weak synoptic forcings are prevalent, nudging only large scales is not expected to considerably improve precipitation forecasts, although extending nudging to the boundary layer may lead to some improvements by mitigating model biases.
@@
Figure 16: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 hr of integration) with (a) equitable threat score, (b) frequency bias, (c) probability of detection, and (d) false alarm ratio. Results from the 60 GDPS-CTL (blue) and GDPS-SN (red) cases of boreal winter are compared to ground observations over North America used by CaPA. Red and blue line markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. Significance is computed by bootstrapping 3-day data blocks consisting of forecast-observation pairs from all stations of a given region. No marker at a threshold level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
@@ 6.5 Tropical cyclone evaluation
Evaluation of the impact of spectral nudging on the tropical cyclone (TC) forecast guidance is performed using the data from the International Best Track Archive for Climate Stewardship (IBTrACS; Knapp et al. 2010). For the results presented in this section, GDPS TC forecasts are matched with observations, by excluding from consideration any storm that is not detected by both GDPS and IBTrACS systems.
Spectral nudging leads to an overall reduction in position error in the predicted TC trajectories (Fig. 17a). Notably, GDPS-CTL’s tendency of predicting storms that move too slowly is improved with GDPS-SN, especially at lead times beyond day 5 (Fig. 17b). There is also an indication of improvement with respect to GDPS-CTL’s tendency of predicting TCs that veer too much to the right from their observed tracks (Fig. 17c). Conversely, there is little or no significant impact of spectral nudging on the TC intensity, as measured by the minimum central pressure (not shown) and maximum sustained wind speed (Figs. 17d). Both GDPS-CTL and GDPS-SN have a similar weak-intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al. (2024), for more details). Severe TC weak-intensity biases are typical of state-of-the-art deterministic AI-based forecast models (e.g., Bouallègue et al. (2024)) due to the smoothing effect caused by learning to minimize the MSE. However, AI-based models, including GC, are generally capable of predicting the TC trajectories with enhanced accuracy (Lam et al., 2022). Therefore, the improvements in the predicted trajectories by GDPS-SN show that the proposed hybrid system is capable of leveraging the well-documented strength of the AI-based models, without sacrificing TC intensity. Furthermore, these results suggest that the weak-intensity bias for TCs could be reduced in future by focusing on improving the physics-based NWP component of the hybrid system.
Figure 17: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East and West Pacific) for 62 cases of boreal summer 2022: (a) mean position error, (b) mean along-track position error, (c) mean cross-track position error, and (d) mean maximum wind speed error. Negative along-track errors imply cyclones move too slowly and positive cross-track errors imply cyclones veer to much to the right with respect to the observed trajectory. Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile. The number of cases for different forecast lead times are indicated in 17c.

### v3

@@ 5.2 Verification against radiosondes
Comparison of forecasts at full resolution against the global radiosonde observations clearly indicates that GDPS-SN forecasts are closer to the observations than GDPS-CTL forecasts across variables, lead times, and levels for both seasons (Fig. 10). The improvements grow gradually over the first 72 hr and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GraphCast in Section 3. This demonstrates that large-scale spectral nudging, despite being only applied to free-tropospheric winds and temperature, is sufficient to impart considerable improvements to the unconstrained upper- and lower-level layers. Compared to boreal winter (Fig. 10a), the improvements from GDPS-SN are generally reduced during boreal summer (Fig. 10b), which is consistent with the relative performance of the component systems (not shown).
@@
Regionally, GDPS-SN has smaller RMSE than GDPS-CTL in every subdomain except over Asia in the mid- and upper-levels during boreal summer (Figs. 11b and d), which is consistent with the degradations noted for GraphCast large scales in this region (Fig. 5d). Unlike GraphCast, the forecast improvements of GDPS-SN over GDPS-CTL, when measured against radiosonde observations, are comparable at both full resolution (left column of Fig. 11) and for large scales only (right column of Fig. 11, compare Fig. 5), showing only a small degradation under the low-pass filter. This implies that the fine scales predicted by GDPS-SN benefit from fractional improvements at large scales, which is supported by the improvements in fine-scale spectral coherence with GDPS-SN over GDPS-CTL, particularly beyond day 3 (Fig. 7c-d).
@@
Figure 10: Same as in Fig. 4, except that upward-pointing red triangles indicate a reduction of RMSE by GDPS-SN with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite. Only verification using forecasts at full resolution is shown.
@@
Figure 11: Same as in Fig. 5, except that red shadings indicate a reduction of RMSE by GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
@@
Figure 12: Same as in Fig. 4, except that upward-pointing red triangles indicate a reduction of RMSE by GDPS-SN with respect to GraphCast, whereas downward-pointing blue triangles indicate the opposite. Only verification using forecasts at low resolution is shown, but for both seasons.
@@
Forecasts from GDPS-SN must also be compared to those from GraphCast to ensure that the skill of the AI-based system is not lost during hybridization. For the sake of fair comparison, this is only done at low resolution (section 3b2). Predictions from GDPS-SN are better than GraphCast in the upper- and mid-level layers, except for dewpoint depression (Fig. 12). Conversely, GraphCast forecasts are closer to observations at lower levels, except for geopotential height. The dewpoint depression degradations suggest that further gains could be achieved by spectrally nudging the humidity field. However, this will only be viable once a fine-tuned version of GraphCast becomes available because of systematic differences between the moisture state estimates of ERA5 and the initializing GDPS analyses. Fine-tuning will bring the mean moisture state of GraphCast inferences closer to GDPS analyses and is expected to reduce errors associated with any adverse response from the GEM physics suite. The lower-level results also suggest that extending spectral nudging to the boundary layer could be beneficial. However, this would require an AI model trained on GEM’s terrain-following vertical coordinate with more vertical levels in the boundary layer, fine-tuned to emulate the GDPS analyses for improved consistency with the GEM model’s lower-boundary forcings.
@@
The small but growing GDPS-SN improvements over GraphCast in the mid-level layer for spectrally-nudged temperature and zonal wind are somewhat surprising (Fig. 12). The nudging relaxation time of $\tau=12$ hr implies that the large scales in GDPS-SN are not too tightly constrained, which is visible in Fig. 8. This may allow for some large-scale improvements through upscale propagation of information from an improved fine-scale representation. However, this choice of $\tau$ also leads to a small averaging effect caused by the blending of two forecasts. To a lesser extent, smoothing attributable to this effect may also have contributed to the noted improvements.
@@
Finally, the regional comparison of GDPS-SN and GraphCast forecasts at low resolution (Fig. 13) shows that the prediction accuracy of GDPS-SN is typically slightly better than or comparable to GraphCast, with the tropical boundary layer being one of the exceptions. Fig. 5 reveals that, compared to GraphCast, GDPS-CTL large scales have substantially reduced accuracy in the tropics, which is slightly improved with GDPS-SN (Fig. 11). To leverage the full potential of GraphCast in the tropics it will likely be necessary to extend nudging in the boundary layer, once fine-tuning to GDPS analyses is complete.
@@
Figure 13: Same as in Fig. 5, except that red shadings indicate a reduction of RMSE by GDPS-SN with respect to GraphCast, whereas blue shadings indicate the opposite. Only verification using forecasts at low resolution is shown, but for both seasons.
@@ 5.3 Verification against ECMWF analyses
The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF analyses – in terms of $f_{RMSE}$ – leads to conclusions that are similar to those of the radiosonde assessment: general improvements for all variables, lead times, and pressure levels (not shown). Figure 14 provides an example of changes in anomaly correlation coefficient (ACC). For 500-hPa geopotential heights, the ACC improvements with GDPS-SN (relative to GDPS-CTL) in day-7 forecasts over the Northern Hemisphere region are equivalent to a 11-hr increase in predictability during boreal winter and an 8-hr increase in summer (Figs. 14a,d). The corresponding ACC improvements over the Southern Hemisphere region are about 3 hours during boreal summer and 13 hours during boreal winter (Figs. 14c,f). There are also notable improvements in zonal wind ACC in the Tropics that are equivalent to 26 hours or more (Figs. 14b,e). The strength of GraphCast in improving the ACC around day 5 (and beyond) is thus well leveraged by the hybrid system. GraphCast has not been included in this figure because its ACC is strongly influenced by the excessive fine-scale smoothing. To fairly compare GraphCast against the two versions of GDPS, it is necessary to do so at low resolution (section 3b); however, based on spectral coherence presented in Figs. 7c–d, low-resolution ACC of GraphCast is expected to be comparable to GDPS-SN.
@@
Figure 14: Anomaly correlation coefficient (ACC) against the ECMWF analyses with GDPS-CTL (blue) and GDPS-SN (red) for (a,d) northern hemisphere 500-hPa geopotential height, (b,e) tropical 850-hPa zonal wind, and (c,f) southern hemisphere 500-hPa geopotential height. Results are presented for 60 cases of boreal winter 2022 (left column: a–c) and 62 cases of boreal summer 2022 (right column: d–f). ACC is computed using climatology of ERA-Interim reanalyses. Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. No marker for a forecast lead time implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 95th percentile. Grey arrows with printed numbers depict gains in predictability in terms of forecast hours.
5.4 Evaluation against surface observations
@@
For near-surface evaluation, combined SYNOP, METAR, and Surface Weather and marine OBservations (SWOB; available only over Canada and distributed by ECCC) data were used for surface pressure, screen-level temperature and dewpoint temperature, and anemometer-level wind speed. For precipitation, ground observations of 24-hr accumulation, subject to collection and quality control by the CAnadian Precipitation Analysis system (CaPA; Lespinas et al. 2015), are used. Any data originating from stations with an altitude difference larger than 100 m with respect to GDPS orography is excluded, whereas any observed or forecast wind speed below 1.5 $\text{m s}^{-1}$ is set to 0.
@@
Figure 15: Heatmap of the changes in the forecast quality index (Eq. 3) against combined SYNOP, SWOB, and METAR observations for surface pressure ($P_{s}$), screen-level temperature ($T$), dewpoint temperature ($T_{d}$), and anemometer-level wind speed ($W$) in various geographical domains for (a) 60 cases of boreal winter 2022 and (b) 62 cases of boreal summer 2022. Results are computed at full resolution. Red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
@@
Figure 16: Q-Q plots of day-5 GraphCast (black), GDPS-CTL (blue), and GDPS-SN (red) forecasts against combined SYNOP, SWOB, and METAR for: (a, b) anemometer-level wind speeds, and (c, d) screen-level temperatures over North America for (a, c) 30 cases of boreal winter valid at (a) 00 and (c) 12 UTC, and (b, d) 31 cases of boreal summer valid 00 UTC. Local standard time over North America is mostly between UTC-5 and UTC-8 h. Results correspond to full resolution of the different models. Diamond symbols denote selected percentiles with their values printed. Light shadings indicate the 5%–95% confidence interval for the inverse cumulative distribution functions, based on the Kolmogorov-Smirnov statistic.
@@
Guidance from GDPS-SN is generally improved over GDPS-CTL for all variables and regions (Fig. 15). The improvements are, however, smaller than in the upper air (Fig. 11). Achieving considerable improvements near the surface appears to be difficult without introducing spectral nudging in the boundary layer. However, spectral nudging in the mid-level layer does have a considerable positive impact on the surface pressure via redistribution of upper-air mass. The improvements during boreal summer are reduced, consistent with the upper-air verification discussed above.
@@
The ability of the GDPS-SN to represent high-impact weather (high wind speeds and extreme temperatures) is not significantly different from the GDPS-CTL (Fig. 16). Both are substantially better than GraphCast, which is unable to predict the tails of the climatological distributions. For example, the distribution of late-afternoon wind speeds over North America provided by GraphCast shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -3 $\text{m s}^{-1}$ bias near ${}^{\text{th}}$ percentile (Figs. 16 a, b). Although GraphCast does not show any discernible bias for low temperature percentiles over North America in winter (Fig. 16 c), it suffers from a 2$\sim$4 °$\degree$°C cold bias at mid-to-high temperature percentiles in summer (Fig. 16 d), a problem that does not affect GDPS-SN. At high temperature percentiles, GraphCast’s poor performance with respect to the extremes can likely be explained by its lack of well-resolved fine scales (Iles et al., 2020) combined with inconsistencies in the surface forcing between ERA5 and GDPS analyses. As the spectral nudging configuration presented in this study only targets synoptic scales, the GEM component of the hybrid system is able to fill in small scales associated with local forcings. GDPS-SN is, therefore, not adversely affected by GraphCast’s limitations regarding the extremes.
@@
Spectral nudging also significantly improves boreal winter precipitation guidance over North America primarily through a reduction in false alarm (Fig. 17). These improvements are likely the direct result of an improved representation of the synoptic-scale flow in strongly forced winter conditions. The impact of spectral nudging on summer precipitation over North America is, however, negligible (not shown). During summer, when weak synoptic forcings are prevalent, nudging only large scales is not expected to considerably improve precipitation forecasts, although extending nudging to the boundary layer may lead to some improvements by mitigating model biases.
@@
Figure 17: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 hr of integration) with (a) equitable threat score (ETS), (b) frequency bias index (FBI), (c) probability of detection (POD), and (d) false alarm ratio (FAR). Results from the 60 GDPS-CTL (blue) and GDPS-SN (red) cases of boreal winter are compared to ground observations over North America used by CaPA. Red and blue line markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. Significance is computed by bootstrapping 3-day data blocks consisting of forecast-observation pairs from all stations. No marker at a threshold level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
@@ 5.5 Tropical cyclone evaluation
Evaluation of the impact of spectral nudging on the tropical cyclone (TC) guidance is performed using the data from the International Best Track Archive for Climate Stewardship (IBTrACS; Knapp et al. 2010) following the methodology described by McTaggart-Cowan et al. (2024).
AI-based models, including GraphCast, are generally capable of predicting the TC trajectories with enhanced accuracy (Lam et al., 2022). Results presented in Fig. 18a show that spectral nudging allows GDPS-SN to leverage GraphCast’s enhanced TC steering accuracy and leads to an overall reduction in position error in the predicted TC trajectories. Notably, GDPS-CTL’s tendency of predicting storms that move too slowly – indicated by negative along-track error – is improved with GDPS-SN, especially at lead times beyond day 5 (Fig. 18b). There is also an indication of improvement with respect to GDPS-CTL’s tendency of predicting TCs that veer too much to the right from their observed tracks, which is indicated by positive cross-track error (Fig. 18c). Conversely, there is little or no significant impact of spectral nudging on TC intensity, as measured by the maximum sustained wind speed (Figs. 18d). Although both GDPS-CTL and GDPS-SN have a similar weak-intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al. (2024), for more details), both versions of GDPS nevertheless predict higher intensity TCs compared to GraphCast. Severe TC weak-intensity biases are typical of state-of-the-art deterministic AI-based forecast models due to the double-penalty effect caused by learning to minimize the MSE (Bouallègue et al., 2024). Combined with the small scales of the TC vortex that make direct nudging of the pertinent scales infeasible, this means that the potential for improvements in intensity prediction lies primarily in the physics-based NWP component of the hybrid system.
Figure 18: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East, and West Pacific) for 62 cases of boreal summer 2022: (a) mean absolute position error, (b) mean along-track position error, (c) mean cross-track position error, and (d) mean maximum wind speed error. Performance of GraphCast (dashed black) is presented for reference. Negative along-track errors imply cyclones move too slowly and positive cross-track errors imply cyclones veer too much to the right with respect to the observed trajectory. Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. No marker at a lead time implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile. The number of cases for different forecast lead times are indicated in 18c.